Gilitch Company has variable costs of $600,000. They are planning to sell 200,000 pairs of mittens...

  1. Gilitch Company has variable costs of $600,000. They are planning to sell 200,000 pairs of mittens for $4 a pair. If Ilitch breaks even at this level of sales what are the fixed costs?
  1. Zark and Sons Ltd. makes travel bags that dissolve in water and has variable costs of $42 per unit and expects fixed costs to total $320,000. The travel bags sell for $56 each. Calculate:

  1. The Break Even Point in Dollars
  1. The Margin of Safety and Margin of Safety ratio assuming actual sales total $1,382,400
  1. The sales dollars required to earn net income of $410,000
Answer #1

Requirement 1

At Break-even point the fixed cost is equal to Contribution margin.

Fixed cost = $200000

Working

Sales revenue $ 800,000
Variable costs $ 600,000
Contribution margin $ 200,000
Fixed cost $ 200,000
Net operating income $ 0

Requirement 2

a---$1,280,000

b---$102,400 and 7%

c---$2,920,000

Working

A Unit Sale price $                    56.00
B Unit Variable cost $                    42.00
C=A - B Contribution margin per unit $                    14.00
D Fixed expenses $ 320,000
E = D/C Units to be sold to breakeven $ 22,857
F = E x A Break Even in dollar sales [Answer 'a'] $ 1,280,000
G Total Sales $ 1,382,400
H = G - F Margin of safety Sales $ 102,400
I = H/G Margin of safety % [Answer 'b'] 7%
Fixed cost+ Profit $ 730,000
Units to be sold to breakeven $ 52,143
Break Even in dollar sales [Answer 'c'] $ 2,920,000